OpenStudy (anonymous):

what's the domain of y=log(x) (assuming that log is base 0)

OpenStudy (noelgreco):

Base 0?

OpenStudy (anonymous):

\[y = \log_{0}(x)\]

OpenStudy (anonymous):

should the domain be x = {0}

OpenStudy (noelgreco):

Ask the question: "To what power do I have to raise zero to get x?"

OpenStudy (anonymous):

That doesn't find the domain.

OpenStudy (cwrw238):

0 ^ n = 0

OpenStudy (anonymous):

Yes. \[\log_0(x):0\to\mathbb{C}\backslash\{0\}\]

OpenStudy (anonymous):

You can raise 0 to the anything, but not 0, or negative.

OpenStudy (anonymous):

why can't you raise 0 to 0?

OpenStudy (anonymous):

Sh*ite: \(\mathbb{C^+}\)

OpenStudy (anonymous):

@LolWolf what does log0(x):0→C∖{0}mean?

OpenStudy (anonymous):

\[0^0 \rightarrow \text{doesn't exist}, 0^{-n} \rightarrow \text{ divide by 0}\]

OpenStudy (anonymous):

It means that, \(\log_0(x)\) will 'map' (has a domain of... roughly speaking) 0 to all of the positive complex numbers, excluding zero.

OpenStudy (anonymous):

Are we doing complex exponents here?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

Well, the reals are part of the complex, so you can just say the same thing, in this case at least, for the reals (without including the complex numbers)

OpenStudy (anonymous):

Then any real number other than 0 or negatives.

OpenStudy (anonymous):

@LolWolf , why DOES maths break at 0^0?

OpenStudy (anonymous):

math is so complicated :S

OpenStudy (anonymous):

Because we can show \(n^0\) to mean \[ \frac{n}{n} \text{ so, it'd be } \frac{0}{0}\text{which is undefined.} \]

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

Because if it does, surely because of:\[a^n=1a^n=a^0a^n=a^{n+0}\], so if a=0, that means that you can't raise 0 to the anythingth power

OpenStudy (anonymous):

I know I'm wrong, but which of my equalities is the fallacious one?

OpenStudy (anonymous):

@henpen : \[ 0^0=1 \]

OpenStudy (anonymous):

Oh, yes- I was writing as you were explaining the previous definition

OpenStudy (anonymous):

so to conclude, if i were to give the domain of \[y = \log_{a}(x)\] it would be a > 0, and x = all real numbers

OpenStudy (anonymous):

sorry, x > 0

OpenStudy (anonymous):

It'd be: \[ a>0, a\neq1\\ x\in\mathbb{R}^+ \]

OpenStudy (anonymous):

Yeah, @artofspeed , better, but make sure \(a \neq 1\)